A centrifuge in the shape of a uniform cylinder of radius 20 cm and mass 10 kg cylinder is spinning about its axis at a speed of 50,000 rpm. Determine the torque that must be applied in order to bring the centrifuge to a stop in 20 s. Calculate the power dissipated.
1
Expert's answer
2020-02-24T11:02:40-0500

Torque:


"τ=Iα=I\frac{\omega}{t}=0.5mr^2\frac{\omega}{t}"

"τ=0.5(10)(0.2)^2\frac{2π(50000)}{20(60)}=52.36\ Nm"

Power:


"P=Tω=52.36\frac{2π(50000)}{20(60)}=274000 W=274 kW"